Circular No. 8934

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2491 CYGNI S. Nakano, Sumoto, Japan, reports the discovery by Koichi Nishiyama

(Kurume, Fukuoka-ken, Japan) and Fujio Kabashima (Miyaki-cho, Sagaken, Japan) of a possible nova (mag 7.7) on two 20-s unfiltered CCD frames (limiting mag ∼ 12.5) taken on Apr. 10.728 UT using a 105-mm-f.l. f/5.6 camera lens; their unfiltered CCD images taken on Apr. 10.787 and 11.668 with a 0.40-m reflector yields α = 19^{h}43^{m}01^{s}.96, δ = +32^{o}19′13′′.8 (equinox 2000.0), and mag 7.7 and 7.1 for the variable; nothing is visible at this position on their past survey frames taken on Apr. 3.717 (limiting mag 12.5) and 7.727 (limiting mag 12.7). Following posting on the Central Bureau’s unconfirmed-objects webpage, J. Beize (Beijing, China) reports the apparent independent discovery of this possible nova at mag ≈ 8.0 by Z.-w. Jin and X. Gao on several 60-s survey images (limiting mag ∼ 15.0) taken by Gao on Apr. 10.831 with a Canon EOS-350D camera (+ 7-cm-aperture, 200-mm-f.l. f/2.8 lens) at Xingming Observatory, Mt. Nanshan, with Beize measuring position end figures 02^{s}.00, 10′′.1 for the variable; nothing was visible at this position on Xingming images taken on Apr. 8.831 (limiting mag ∼ 14). H. Yamaoka, Kyushu University, writes that K. Haseda (Toyohashi, Aichi, Japan) finds nothing brighter than mag 12.3 at the reported position for the variable on his patrol images taken on Apr. 4.774 with a Canon EOS-5D digital camera + 120-mm-f.l. lens. E. Guido and G. Sostero, Remanzacco, Italy, report that their confirming CCD images taken remotely with a 0.25-m reflector near Mayhill, NM, U.S.A., on Apr. 11 show the apparent nova at position end figures 01^{s}.98, 13′′.5 (UCAC-2 catalogue reference stars). Sostero and Guido add that comparison with a Palomar Oschin Schmidt telescope plate obtained on 1995 Aug. 3, shows the presence at this position of an apparent star with red mag ∼ 18; they also provide the following magnitudes (uncertainty ± 0.05 mag in each color): Apr. 11.36, B = 8.06, V = 7.54, R = 7.04; 11.46, 8.20, 7.77, 7.15. G. Klingenberg, Mo i Rana, Norway, reports the following magnitudes for the variable from CCD images taken with a 25-cm reflector near Mayhill: Apr. 11.333, V = 7.63; 11.335, B = 8.06; 11.440, V = 7.74; 11.443, B = 8.15. Nakano adds that an unfiltered CCD taken by K. Kadota (Ageo, Saitama-ken, Japan, 0.25-m reflector) on Apr. 11.741 yields mag 7.4 and position end figures 01^{s}.98, 14′′.1; Kadota notes that a USNO-B1.0-catalogue star has position end figures 02^{s}.039, 13′′.84 and red magnitudes 15.5 and 16.3 (different epochs evidently from Palomar Sky Survey plates).

N. N. Samus, Institute of Astronomy, Moscow, informs us that this object is being assigned the variable-star designation V2491 Cyg.

2008 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green